OpenStudy (anonymous):

ALGEBRA 2 HELP PLEASE QUESTION WILL BE IN THE DISCUSSION

OpenStudy (anonymous):

The function p(x) = 2x +1 determines how many pizzas need to be purchased for an after school meeting, where x is the number of students at the meeting. The club leader uses r(p(x)) to find the amount of money to bring for the pizza purchase. The function r(x) = 4x − 3. Solve for how much money to bring when there are 9 students in the meeting

OpenStudy (anonymous):

@perl

OpenStudy (perl):

so you need to find r(p(9))

OpenStudy (anonymous):

I have been on this question for 10 minutes. I honestly dont know how to work this out. can you please walk me through it

OpenStudy (perl):

lets find an expression for r(p(x))

OpenStudy (anonymous):

p(9) = 2(9) + 1 = 19 p(x)=19 r(p(9)) = r(19) = 4(19) - 3 = 73

OpenStudy (anonymous):

Thank you @moralmer000

OpenStudy (anonymous):

you are very welcome

OpenStudy (anonymous):

On the graph of the equation 5x − 2y = 20, what is the value of the y-intercept?

OpenStudy (anonymous):

@moralmer000 @perl

OpenStudy (anonymous):

y=-5/2x-10 y= mx+b M= slope b= y-intercept

OpenStudy (anonymous):

so y-intercept = -10

OpenStudy (anonymous):

What is the equation of a vertical line passing through the point (−4, 7)

OpenStudy (anonymous):

@moralmer000